OpenStudy (anonymous):

differentiate the function g(r)= r^2(ln(2r+1))

OpenStudy (anonymous):

product rule plus chain rule \[(fg)'=f'g+g'f\] with \[f(r)=r^2, f'(r)=2r,g(r)=\ln(2r+1), g'(r)=\frac{2}{2r+1}\]